Define h(0,0) in such a way that h is continuous at origin here n=5

h(u,v)=ln⁡((nu^2-u^2 v^2+nv^2)/(u^2+v^2 ))


1
Expert's answer
2021-05-07T09:49:04-0400

Checking continuity at (0,0)

"\lim _{x\to \:0}(f(x))"


"\\lim\\limits_{u,v \\to (0,0)}h(u,v)=\\lim\\limits_{u,v \\to (0,0)} (\\frac{ln\u2061((nu^2-u^2 v^2+nv^2)}{(u^2+v^2 )})"


Left hand side

"\\lim\\limits_{u,v \\to (0,0)} (\\frac{ln\u2061((5u^2-u^2 v^2+5v^2)}{(u^2+v^2 )})=-\\infin"


Right-hand side

"\\lim\\limits_{u,v \\to (0,0)} (\\frac{ln\u2061((5u^2-u^2 v^2+5v^2)}{(u^2+v^2 )})=-\\infin"

h is not contiuous at (0,0)